# An analysis of debris-induced nonlinear phenomena in a positive ion-negative ion (PINI) plasma

### v1

@@ Abstract
In this work, an analysis of nonlinear waves and phenomena induced by an external charged debris in a positive ion-negative ion (PINI) plasma is presented. The process of formation of different nonlinear structures is examined theoretically through a forced Koretweg-de Vries (fKdV) equation, which is also verified with a multi-fluid flux-corrected transport (mFCT) simulation. The different processes which are responsible for different nonlinear structures excited by differently charged external debris are pointed out. This work also tries to point out the similarities in different nonlinear structures by an external charged debris (this work) and the underlying processes and those produced by a positive ion beam (other works) in a PINI plasma through beam-plasma interaction.
@@ I Introduction
Nonlinear waves ubiquitously drive the dynamics of nature’s most intricate systems, shaping complex phenomena from plasma instabilities to oceanic rogue waves in ways both beautiful and unpredictable. Plasma shocks and solitons remain prime examples of such structures that governs the energy transport and wave dynamics across laboratory and astrophysical environments. Solitons or solitary waves (SWs) can be considered as the localized compressions or rarefactions of plasma densities arising out of the interplay between nonlinearity and dispersion in a medium. On the contrary, shock waves (SHWs) are usually associated with some external dissipation mechanism such as kinematic viscosity, Landau damping, and collision among ions and neutrals along with inherent nonlinearity, dispersion, and weak dissipation. During a nonlinear ion-acoustic wave (IAW) propagation in a dissipative plasma, the leading edge of a propagating wave gets steepened as dissipation dominates over dispersion and a shock front is formed. In contrast to this, in a collision-less dispersive fluid, SHWs excited without dissipation often have an oscillatory signature and are known as dispersive shock waves (DSWs). Nevertheless, the nature of these nonlinear structures is largely determined by the background plasma conditions, constituents and the physical parameters associated.
@@
The characteristics of IA solitons or SHWs excited in an usual $e$-$i$ plasma are significantly different than that of a multi-ion plasma or a positive ion-negative ion (PINI) plasma. In fact, an $e$-$i$ plasma can only support compressive solitons with positive electrostatic potential. However, with an addition of even a small amount of one or more extra ion species to the plasma, making it multicomponent, drastically changes the dynamics of the plasma. It is observed that in a multicomponent plasma, the IA wave propagation is mainly influenced by the lighter ions and changes in their concentrations and temperatures may lead to substantial modification of the wave structure via the plasma potential [1, 2, 3]. If the added ion species is negatively ionized, then after a critical concentration of these negative ions, the plasma potential can even change its polarity to form a negative or rarefactive soliton [2, 3]. When an ion beam is injected into such a plasma, beam-plasma interaction makes the evolution much more complex in terms of nonlinear wave structures compared to a simple plasma [3, 4, 5]. For this reason, investigations on the plasmas with negative ions as one of the constituents has always been a topic of interest. Plasmas in astrophysical and space environments such as cometary tails, planetary ionospheres, and interstellar and molecular clouds usually are inherently multi-species with two or more than two types of ions and offer a rich platform for studying a variety of wave motions with subtle intricacies [6, 7]. These multi-species plasmas with negative ions are also easily produced in a laboratory and are well studied owing to their vital role in major scientific and technological applications such as plasma processing or fusion plasmas. We believe that an investigation of nonlinear wave propagation in a PINI plasma embedded with an external charged debris is definitely a topic worth studying.
I.1 Can other phenomena mimic debris-induced nonlinearities?
@@
Effect of presence of external debris in a flowing plasma has definitely gathered attention in recent years, especially as awareness and knowledge about interaction of space debris in low-earth orbits with ionospheric plasmas have come to the forefront of space research [8, 9, 10, 11, 12, 13, 14, 15]. Though science of detection of space debris through debris-induced plasma interaction is still quite speculative, we are definitely learning more and more about it, as research intensifies in this area. In this context, we would like to raise a fundamental question about the similarities of the dynamics and structures of nonlinear waves in a plasma in general, to those induced by external charged debris. We do hope that we shall be able clear some issues about this question through the present work.
@@
In many experiments involving beam-plasma interactions, the beam is produced in situ, which leaves the plasma quasi-neutral. Different nonlinear structures produced due to these interactions have interesting dynamical signatures. In such cases, if the beam velocity $v_{b}$ is $\ll v_{s}$, the effective sound velocity, the propagating ion-acoustic wave (IAW) does not see the Debye-scale disturbances created by the beam. This is analogous to what is known as ‘plasma approximation’, which requires $(k\lambda_{D})^{2}\ll 1$, where $k$ is the wave number of the IAW and $\lambda_{D}$ is the electron Debye shielding length. We argue that so long as $v_{b}$ remains considerably smaller than $v_{s}$, the plasma quasi-neutrality is maintained. However as $v_{b}\sim v_{s}$, the beam-plasma dynamics start getting affected by the IAW. And when $v_{b}\gg v_{s}$, the situation becomes quite different and the beam dynamics will be completely detached from the ion-acoustic dynamics. If we now consider the ion-acoustic time scale $\tau_{{\rm IA}}$, it comes out to be $\sim 0.01-0.5\,{\rm milli}\,{\rm second}$ for a typical laboratory device. On the other hand, the electron-ion collision time $\tau_{ei}$ is quite large $\sim 5\,{\rm milli}\,{\rm second}$ for such parameters. As a result, when $v_{b}\gg v_{s}$, there is no way that thermalization can occur between the beam and the background plasma and the background plasma should see the beam as an external charged perturbation (debris) and the beam-plasma dynamics should closely mimic the dynamics observed in plasma-debris interaction. In subsequent sections, we show that this is indeed true and can be proven quite reasonably. We also see similarities among the nonlinear structures produced by externally induced perturbation in plasma to that of debris-induced perturbation. In this regard debris-induced perturbations can be thought to be quite commonplace in many experimental arrangements.
@@
In this work, we explore different nonlinear structures excited by a moving charged debris in a multicomponent PINI plasma and also look at the similarities of their characteristics with experimental results. The complete plasma system is modeled using a forced Korteweg–de Vries (fKdV) equation and results from theoretical analysis were verified using a 1-D multi-fluid flux-corrected transport (mFCT) simulation [13]. The paper is organized as follows. The detailed plasma model and the governing equations are described in Sec. II. Sec. III contains the description of fKdV dynamics and interpretation of its results. In Sec. IV results from FCT simulation were presented, Sec. V briefly addresses the variation of nonlinearity with debris velocity and in Sec. VI, we conclude.
@@ II Plasma model and governing equations
The model we are going to consider is a 1-D warm positive ion-negative ion (PINI) plasma with an external charge debris embedded into it. The relevant equations are the continuity and momentum equations for positive and negative ions. The equations are closed by the Poisson equation. The electrons are considered to be Boltzmannian. The equations are given by
@@
$n_{-}\frac{dv_{-}}{dt}$ $=$ $-\frac{\sigma}{\mu}\frac{\partial n_{-}}{\partial x}+\frac{n_{+}}{\mu}\frac{\partial\phi}{\partial x},$ (3)
@@
where $(n_{\pm},v_{\pm})$ are respectively positive and negative ion densities and velocities, $\phi$ is the plasma potential. The quantity $\sigma=T_{i}/T_{e}$ is the ratio of ion temperature to that of the electrons. Here, we have taken the temperatures of positive and negative ions to be equal $T_{i}=T_{+}=T_{-}$. The mass ratio of negative to positive ions is denoted by $\mu=m_{-}/m_{+}$ and $\delta=n_{+0}/n_{e0}$ is the ratio of the equilibrium positive ion density to the equilibrium electron density. In order to be able to compare our analytical results with existing laboratory experiments, we have assumed $m_{+}>m_{-}$, which is mostly the case in case of experiments [2, 3, 4, 5, 16]. In the above equations, the densities are normalized by their respective equilibrium values and potential is normalized by $(T_{e}/e)$. Time is normalized by $\omega_{i}^{-1}$, with $\omega_{i}=\sqrt{n_{e0}e^{2}/(m_{+}\epsilon_{0})}\equiv\delta^{-1/2}\omega_{p+}\equiv(\delta-1)^{-1/2}\omega_{p-}$, where $\omega_{p\pm}$ are the respective positive ion and negative ion plasma frequencies. While length is normalized by electron Debye length, velocities are normalized by the ion sound speed $c_{s}=\sqrt{T_{e}/m_{+}}$, as determined by the mass of the heaviest species. The quantity $\rho_{{\rm ext}}\equiv\rho_{{\rm ext}}(x-v_{d}t)$ is the normalized charge density of the external debris with $v_{d}$ as the normalized velocity of the external debris. The overall charge neutrality of the PINI plasma without external debris is ensured through the relation
@@ II.1 The morphology of debris-induced nonlinearities
Before we investigate the nonlinear waves and structures induced by a charged debris in a PINI plasma, it is of relevance to discuss a bit about the fundamental physics of these nonlinearities in an $e$-$i$ plasma. We now know that the response of an $e$-$i$ plasma to an embedded charged debris very well depends on the nature of the charge of the debris [13]. It has been only recently shown that the formation of bright pinned solitons in the ion-acoustic regime due to the presence of a negatively charged debris is basically a manifestation of the trapped ions in phase-space vortices formed due to ion-ion counter streaming instability (IICSI) [15]. Besides, proper spatial resolution of the pinned solitons requires a certain relative velocity between the debris and the plasma, though higher relative velocities may cause the pinned solitons to disappear [10, 15]. In contrast, a positively charged debris creates an ion hole, exciting an ion-acoustic wave (IAW), which propagates away from the site of the debris. However, when there is a relative motion between the debris and the plasma, the IAW becomes a dispersive shock wave (DSW) in the precursor region [13].
Following the above explanation, we expect that the response of a PINI plasma to external charged debris should also be characteristically different for positively and negatively charged debris. We further expect that for $\delta\sim 1$, the response should be similar to that of an $e$-$i$ plasma.
@@ III Forced-KdV dynamics
where $\varepsilon$ is the expansion parameter, which is $\ll 1$. We now use the stretched variables $\xi=\varepsilon^{1/2}(x-Vt)$ and $\tau=\varepsilon^{3/2}t$ for space and time coordinates, where $V$ is the dimensionless phase velocity of the IAW. Following the usual procedure, by collecting the first order terms, we have the following expressions for the first order variables in terms of $\phi^{(1)}$,
@@
We note that unlike the KdV equation, the fKdV equation, in general, is non-integrable and has to be solved numerically. In what follows, we present the numerical solutions of Eq.(14) for a Gaussian charge density profile for the external debris with periodic boundary conditions.
III.1 Numerical results
We now present the results representing nonlinear structures in a flowing PINI plasma with an embedded external charged debris, as obtained from the numerical solution of the fKdV equation given by Eq.(14). What we shall observe is the appearance of dark pinned solitons as debris velocity becomes supersonic when the debris charge is positive. In what follows, we shall discuss two other experimental cases (i) PINI plasma with a positive ion-beam and (ii) large-amplitude perturbation in a PINI plasma, for which we shall try to draw a parallel between the nonlinear oscillations for a PINI plasma with external charged debris (this study) and these experimental cases. All the figures in the subsequent sections are drawn in the rest frame of the debris. We present our results with respect to two independent parameters – $\delta$ and $v_{d}$ for a fixed $\sigma$.
III.1.1 Perturbation with positively charged external debris

### v3

# Exploring nonlinearities in a positive ion-negative ion (PINI) plasma: can other processes mimic debris-induced effects?

@@ Abstract
In this work, an analysis of nonlinear waves and structures induced by an external charged debris in a positive ion-negative ion (PINI) plasma is presented. The results obtained are compared with findings from available experiments involving PINI plasma. The process of formation of different nonlinear structures is examined theoretically through a forced Koretweg-de Vries (fKdV) equation, which is also verified with a multi-fluid flux-corrected transport simulation code (mFCT). Various processes which are responsible for different nonlinear waves and structures excited by differently charged external debris are pointed out. This work also points out the similarities in different nonlinear structures excited by an external charged debris and the underlying processes (this work) and those observed experimentally through processes that do not involve any external debris.
@@ I Introduction
Nonlinear waves ubiquitously drive the dynamics of nature’s most intricate systems, shaping complex phenomena from plasma instabilities to oceanic rogue waves in ways both beautiful and unpredictable. Plasma solitons and shocks remain prime examples of such structures that govern the energy transport and wave dynamics both in laboratory and astrophysical environments. Solitons or solitary waves can be considered as localized compressions or rarefactions of plasma densities arising out of the interplay between nonlinearity and dispersion. On the contrary, shock waves are usually associated with some dissipation mechanisms such as kinematic viscosity, Landau damping, and collision among ions and neutrals along with inherent nonlinearity, dispersion, and weak dissipation. During the nonlinear evolution of ion-acoustic wave (IAW) in a dissipative plasma, the leading edge of a propagating wave gets steepened as dissipation dominates over dispersion and a shock front is formed. In contrast to this, in a collision-less dispersive fluid, shock waves can form without dissipation and often have an oscillatory signature and are known as dispersive shock waves (DSWs) [1, 2, 3, 4]. Nevertheless, the nature of these nonlinear structures is largely determined by the background plasma conditions, constituents, and the associated physical parameters.
@@
The characteristics of ion-acoustic solitons or shock waves in a usual electron-ion ($e$-$i$) plasma are significantly different from that of a multi-ion plasma such as a positive ion-negative ion (PINI) plasma [5, 6, 7, 8]. In fact, an $e$-$i$ plasma can only support compressive solitons with positive electrostatic potential. However, with an addition of even a small amount of one or more extra ion species to the plasma, drastically changes the dynamics of the plasma. It is observed that in a multicomponent plasma, the IAW propagation is mainly influenced by the lighter ions and changes in their concentrations and temperatures may lead to substantial modification of the wave structure via the plasma potential [9, 10, 11]. If the added ion species is negatively ionized, then after a critical concentration of these negative ions, the plasma potential can even change its polarity to form a negative or rarefactive soliton [10, 11]. When an ion beam is injected into such a plasma, beam-plasma interaction makes the evolution even more complex in terms of nonlinear wave structures compared to a simple plasma [11, 12, 13]. For this reason, investigations on the plasmas with negative ions as one of the constituents has always been a topic of interest. Plasmas in astrophysical and space environments such as cometary tails, planetary ionospheres, and interstellar and molecular clouds usually are inherently multi-species with two or more than two types of ions and offer a rich platform for studying a variety of wave motions with subtle intricacies [14, 15]. These multi-species plasmas with negative ions are also easily produced in a laboratory and are well studied owing to their vital role in major scientific and technological applications such as plasma processing or fusion plasmas [8]. However, the response of a PINI plasma to an external charged debris is a topic yet to be explored fully. Moreover, as we explain in the next subsection, we raise a fundamental question about plasma-debris interaction – can other nonlinear processes mimic debris-induced nonlinearities? As we shall see, under certain circumstances, this is indeed true. So, we believe that an investigation of nonlinear wave propagation in a PINI plasma embedded with an external charged debris is definitely a topic worth studying, which should help us in understandings about the intricacies of the processes leading to formation of these nonlinear structures in multicomponent plasma.
Earlier and recent studies by several authors [16, 17, 18, 19] on phenomena involving debris-induced nonlinearities identified the compressive structures in the precursor and wake regions of debris as well as pinned solitons at the location of the debris. This work shows that rarefactive structures can also be excited by debris, owing to the presence of negative ions. Notably, for the same debris velocity, the morphology of the excited waves in the three regions varies significantly with the concentration of negative ions. The new findings also show certain qualitative similarities with structures observed in other experimental processes suggesting a broader universality in the underlying nonlinear wave excitation mechanisms.
I.1 Can other processes mimic debris-induced nonlinearities?
@@
Effect of presence of external debris in a flowing plasma has definitely gathered attention in recent years, especially as awareness and knowledge about interaction of space debris in low-earth orbits with ionospheric plasmas have come to the forefront of space research [16, 20, 17, 18, 21, 22, 23, 19]. Though science of detection of space debris through debris-induced plasma interaction is still quite speculative, we are definitely learning more and more about it, as research intensifies in this area. In this context, we would like to raise a fundamental question about the similarities of the dynamics and structures of nonlinear waves in a plasma in general, to those induced by external charged debris. We do hope that we shall be able to clear some issues about this question through the present work.
@@
In many experiments involving beam-plasma interactions, the beam is produced in situ, which leaves the plasma quasi-neutral. Different nonlinear structures produced due to these interactions have interesting dynamical signatures. In such cases, if the beam velocity $v_{b}$ is $\ll v_{s}$ ($v_{s}$is the effective sound velocity), the propagating IAW does not see the Debye-scale disturbances created by the beam. This is analogous to what is known as ‘plasma approximation’, which requires $(k\lambda_{D})^{2}\ll 1$ to be applicable, where $k$ is the wave number of the IAW and $\lambda_{D}$ is the electron Debye shielding length. We argue that so long as $v_{b}$ remains considerably smaller than $v_{s}$, the plasma quasi-neutrality is maintained. However as $v_{b}\sim v_{s}$, the beam-plasma dynamics start getting affected by the IAW. And when $v_{b}\gg v_{s}$, the situation becomes quite different and the beam dynamics will be completely detached from the ion-acoustic dynamics. If we now consider the ion-acoustic time scale $\tau_{{\rm IA}}$, it comes out to be $\sim 0.01-0.5\,{\rm milli}\,{\rm second}$ for a typical laboratory device. On the other hand, the electron-ion collision time $\tau_{ei}$ is quite large $\sim 5\,{\rm milli}\,{\rm second}$ for such parameters. As a result, when $v_{b}\gg v_{s}$, there is no way that thermalization can occur between the beam and the background plasma and the background plasma should see the beam as an external charged perturbation (debris) and the beam-plasma dynamics should closely mimic the dynamics observed in plasma-debris interaction. In subsequent sections, we show that this is indeed true and can be proven quite reasonably. We also see similarities among the nonlinear structures produced by externally induced perturbation in plasma to that of debris-induced perturbation. In this regard debris-induced perturbations can be thought to be quite commonplace in many experimental arrangements.
@@
In this work, we explore different nonlinear structures excited by a moving charged debris in a multicomponent PINI plasma and also look at the similarities of their characteristics with different experimental results. The complete plasma system is modeled using a forced Korteweg–de Vries (fKdV) equation and results from theoretical analysis are verified using a 1-D multi-fluid flux-corrected transport simulation code (mFCT) [22]. The paper is organized as follows. The detailed plasma model and the governing equations are described in Section II. Section III contains the description of the fKdV dynamics and interpretation of its results. The results of Section III are verified using FCT simulation in Section IV. In Section V, we briefly address the variation of nonlinearity with debris velocity and in Section VI, we conclude.
@@ II Plasma model and governing equations
The model we are going to consider is a 1-D warm positive ion-negative ion (PINI) dissipation-less plasma with an external charge debris embedded into it. The relevant equations are the continuity and momentum equations for positive and negative ions. The equations are closed by the Poisson equation. The electrons are considered to be Boltzmannian. The equations are given by
@@
$n_{-}\frac{dv_{-}}{dt}$ $=$ $-\frac{\sigma}{\mu}\frac{\partial n_{-}}{\partial x}+\frac{n_{-}}{\mu}\frac{\partial\phi}{\partial x},$ (3)
@@
where $(n_{\pm},v_{\pm})$ are positive and negative ion densities and velocities respectively and $\phi$ is the plasma potential. The quantity $\sigma=T_{i}/T_{e}$ is the ratio of ion temperature to that of the electrons. Here, we have taken the temperatures of positive and negative ions to be equal $T_{i}=T_{+}=T_{-}$. The mass ratio of negative to positive ions is denoted by $\mu=m_{-}/m_{+}$ and $\delta=n_{+0}/n_{e0}$ is the ratio of the equilibrium positive ion density to the equilibrium electron density. In order to be able to compare our analytical results with existing laboratory experiments, we have assumed $m_{+}>m_{-}$, which is mostly the case in case of experiments [10, 11, 12, 13, 24]. In the above equations, the densities are normalized by their respective equilibrium values and potential is normalized by $(T_{e}/e)$ with temperature expressed in energy unit. Time is normalized by $\omega_{i}^{-1}$, with $\omega_{i}=\sqrt{n_{e0}e^{2}/(m_{+}\epsilon_{0})}\equiv\delta^{-1/2}\omega_{p+}\equiv(\delta-1)^{-1/2}\omega_{p-}$, where $\omega_{p\pm}$ are the respective positive ion and negative ion plasma frequencies. While length is normalized by electron Debye length, velocities are normalized by the ion sound speed $c_{s}=\sqrt{T_{e}/m_{+}}$, determined by the mass of the heaviest species. The quantity $\rho_{{\rm ext}}\equiv\rho_{{\rm ext}}(x-v_{d}t)$ is the normalized charge density of the external debris with $v_{d}$ as the normalized velocity of the external debris. The overall charge neutrality of the PINI plasma without external debris is ensured through the relation
@@
An explanation about the ‘dissipation-less’ assumption is given in Section III (just before Section III-A), where it is more appropriate to discuss it with reference to the KdV equation.
@@ II.1 The morphology of debris-induced nonlinearities
Before we investigate the nonlinear waves and structures induced by a charged debris in a PINI plasma, it is of relevance to discuss a bit about the fundamental physics of these nonlinearities in an $e$-$i$ plasma. We now know that the response of an $e$-$i$ plasma to an embedded charged debris very well depends on the nature of the charge of the debris [22]. It has been only recently shown that the formation of bright pinned solitons in the ion-acoustic regime due to the presence of negatively charged debris is basically a manifestation of the trapped ions in phase-space vortices formed due to ion-ion counter streaming instability (IICSI) [19]. Besides, proper spatial resolution of the pinned solitons requires a certain relative velocity between the debris and the plasma, though higher relative velocities may cause the pinned solitons to disappear [17, 19]. In contrast, a positively charged debris creates an ion hole, exciting an ion-acoustic wave (IAW), which propagates away from the site of the debris. However, when there is a relative motion between the debris and the plasma, the IAW becomes a dispersive shock wave (DSW) in the precursor region [22].
Following the above explanation, we expect that the response of a PINI plasma to external charged debris should also be characteristically different for positively and negatively charged debris. We further expect that for $\delta\sim 1$, the response should be similar to that in an $e$-$i$ plasma.
II.2 Mathematical reduction
We note that the above plasma model can be analyzed both mathematically and numerically. Typically, in these class of problems, one uses a scale transformation and reduces it to a pseudo potential (or Sagdeev potential) form [25] through the Poisson equation, which guarantees a soliton solution, though not necessarily a ‘${\rm sech}$’-type solution as one would have obtained through a KdV equation [26]. The other two approaches would be a reductive perturbation analysis to reduce the system to a KdV equation or a nonlinear Schrödinger equation (NLSE) [27]. However, with an external charged debris term, the model is simply not amenable to pseudo-potential analysis and that leaves us with other two options.
Employing the reductive perturbation method with the external charged debris term, we end up with either a forced-KdV (fKdV) or a forced-NLSE equation. At this point, we should emphasize that numerically we are going to a use flux-corrected transport (FCT) method (as mentioned before) to simulate the dynamics, which is very efficient in detecting sharp discontinuities that can result from the debris-induced perturbation. Also the FCT formalism does not employ any approximations and solves the full set of equations, which in principle should provide us with both KdV-type as well as NLSE-type modulated wave solutions. However, the single hump perturbation that we expect from an external charged debris, the numerical solution is most likely to favor KdV-type solution which is also closer to the real-world and observable dynamics. In what follows, we shall therefore analyze the above model mathematically, using the fKdV equation only.
@@ III Forced-KdV dynamics
where $\varepsilon$ is the expansion parameter, which is $\ll 1$. One question related to the order of perturbation for the external charged debris may naturally arise here. In this context, two points to be noted here – that we use the KdV equation for weak nonlinearity only and how strong a perturbation an external charged debris may induce in the bulk plasma. Our assumption of using a second order perturbation for the debris is quite in line with the weak nonlinearity assumption on the basis of which the KdV equation is derived. As far as the strength of perturbation is concerned, we shall see in subsequent sections that the experimentally observed nonlinearities are quite close to our numerical solutions (obtained via FCT simulation) as well as the fKdV solutions, which demonstrates that our assumption of introducing the external charged debris as second order perturbation is indeed consistent. Besides, using a second order perturbation for debris leaves us with the simplest form of fKdV equation.
We now use the stretched variables $\xi=\varepsilon^{1/2}(x-Vt)$ and $\tau=\varepsilon^{3/2}t$ for space and time coordinates, where $V$ is the dimensionless phase velocity of the IAW. Following the usual procedure, by collecting the first order terms, we have the following expressions for the first order variables in terms of $\phi^{(1)}$,
@@
We note that unlike the KdV equation, the fKdV equation, in general, is non-integrable and has to be solved numerically. So, we present the numerical solutions of Eq.(14) for a Gaussian charge density profile for the external debris with periodic boundary conditions. Here, we consider two possible scenarios – positively charged debris (Section III-A) and negatively charged debris (Section III-B). For positively charged debris, we also compare the debris-induced nonlinearity with results from two different experiments, both of which employ different mechanisms to excite nonlinear waves. All the figures in the subsequent sections are drawn in the rest frame of the debris. We present our results with respect to two independent parameters – $\delta$ and $v_{d}$ for a fixed $\sigma$.
As mentioned before, we now provide an explanation about our ‘dissipation-less’ assumption of the plasma model. The dissipation in this case may come in two flavors – through collisions and Landau damping. There are examples of work involving multicomponent plasma (not necessarily PINI plasma) with collisions as well as with Landau damping term [28]. While the inclusion of different collisional terms in the fluid model is quite straight forward [29, 30], the same is not true in the case of Landau damping. However, as shown by VanDam [31], it is indeed possible to include a dissipation-like term arising out of wave-particle resonance (Landau damping) in a KdV-like formalism. In Section III-A.1, while comparing our fKdV solutions with ion-beam-induced nonlinearities, we shall see that the $e$-$i$ collision time $\tau_{ei}\sim 3.5\,{\rm milli}\,{\rm second}$. However, the ion-acoustic time scale, which is the characteristic time scale of this work, is $\tau_{{\rm IA}}\sim 0.03\,{\rm milli}\,{\rm second}$. So, we can probably safely neglect collisional dissipation. Coming back to now Landau damping, VanDam [31] has shown that a dissipation-like term can be incorporated into the KdV equation with the linear Landau damping rate $\gamma=\tilde{\gamma}|k|$ (see Appendix A) where
$\tilde{\gamma}=\frac{\lambda}{\alpha\varepsilon\sqrt{2\pi}}\left[\left(\frac{m_{e}}{m_{+}}\right)^{1/2}+\sigma^{-3/2}\exp\left(-\frac{\lambda^{2}}{2\sigma}\right)\right],$ (17)
with
$\lambda$ $=$ $1+\frac{3}{2}\sigma,$ (18)
$\alpha$ $=$ $2\lambda^{-2}-3\lambda^{-4}\sigma.$ (19)
In the above expression, $k$ is the wavenumber of the associated IAW. The expansion parameter can be taken as $\varepsilon\sim\sigma^{-3/2}e^{-1/(2\sigma)}$ [31]. This expression is written with the assumption that the momentum of the resultant IAW in a PINI plasma is carried primarily by the heavier positive ions. With typical parameters used in this work (see Section III-A.1), we find that the damping timescale $\tau_{\gamma}$ corresponding to $\gamma$ is $\sim 1\,{\rm milli}\,{\rm second}$, which is more than two orders of magnitude larger than $\tau_{{\rm IA}}$. So, owing to the fact that $\tau_{\gamma}\gg\tau_{{\rm IA}}$, we can safely neglect the effect of Landau damping in this work.
III.1 Positively charged external debris